# Deep Inverse Design for High-Level Synthesis

### v2

@@ Abstract
High-level synthesis (HLS) has significantly advanced the automation of digital circuits design, yet the need for expertise and time in pragma tuning remains challenging. Existing solutions for the design space exploration (DSE) adopt either heuristic methods, lacking essential information for further optimization potential, or predictive models, missing sufficient generalization due to the time-consuming nature of HLS and the exponential growth of the design space. To address these challenges, we propose Deep Inverse Design for HLS (DID4444HLS), a novel approach that integrates graph neural networks and generative models. DID4HLS iteratively optimizes hardware designs aimed at compute-intensive algorithms by learning conditional distributions of design features from post-HLS data. Compared to four state-of-the-art DSE baselines, our method achieved an average improvement of $42.8\%$ on average distance to reference set (ADRS) compared to the best-performing baselines across six benchmarks, while demonstrating high robustness and efficiency.
Index terms: ADRS, deep learning, generative model, GNN, HLS, IR, Pareto front
Keywords: deep learning, GAT, HLS, inverse design, post-HLS IR, VAE
@@ I Introduction
Heuristic methods offer rapid implementations to yield improved designs [5, 6, 7, 8, 9, 10]. However, their effectiveness is limited by the lack of a thorough understanding of the behaviors of HLS tools. We need to understand how operations are scheduled in order to estimate design performance, but scheduling information is only available after the HLS process [11]. Recently, graph neural networks (GNNs) have been employed for code analysis to predict synthesis results, demonstrating advantages over the preceding DSE methods [12, 13, 14, 15, 16]. Compilers transform algorithms in HLL into platform-independent intermediate representations (IRs) that encapsulate semantic information [17, 18, 19]. They are subsequently transformed into control and data flow graphs (CDFGs) for GNN training [20, 21]. However, three critical issues hinder the effectiveness of these methods:
$1$) The exponential increase in design space size and the time-consuming nature of HLS pose challenges in acquiring sufficient synthesis data for training high-generalization models [15, 22].
@@
Gradient-based inverse design has been widely adopted to retrieve the nanophotonic device designs with desired quality of results (QoR) from large design spaces [24, 25]. However, the inherent discrete nature of HLS design spaces keeps from applying the inverse design to DSE for HLS. Although several recent approaches approximated discrete spaces with continuous spaces [26, 27], they are not feasible for the inverse design in HLS due to the fact that the pragma effect on the QoR is non-monotonic [13].
To address these challenges, we propose a novel method called Deep Inverse Design for HLS (DID4444HLS). DID4444HLS focuses on choosing the optimal pragma configurations of the compute-intensive algorithms to achieve designs that are close to Pareto points. We construct the design space with the pragmas of pipelining, loop unrolling, array partitioning, and function inlining. The uniqueness of DID4444HLS lies in its strategy to work on a distribution that samples pragma configurations, with constraints that prevent redundant design samples according to the strategies of HLS tools, such as interactions between pipelining and loop unrolling. Initially, DID4444HLS uses a random pragma distribution to explore a wide range of design possibilities. It then iteratively synthesizes batches of designs generated from the updated pragma distributions, resulting in the IRs that are generally represented in CDFG and crucial in deriving their final hardware designs [28, 29]. In this work, we call these IRs as post-HLS IRs (PIRs). These PIRs, along with data on their objectives, are collected into a dataset which is used to train the objective predictors based on Graph Attention Network v2222 (GATv2222) [30]. Furthermore, DID4444HLS adopts Variational Autoencoder (VAE) to learn the distribution of design features that determine design performance. They are extracted from the PIRs by the GATv2222 models. This training is conditioned on the pragma distributions. Through this process, DID4444HLS effectively converges the pragma distribution to approximate Pareto front [3].
@@ II-A Post-HLS Intermediate Representation
The initial HLS phase parses algorithms in HLL into virtual instruction sets called intermediate representations. Subsequent phases entail specifying the order and timing of these instructions under predefined constraints to generate PIRs [28, 29]. PIR is a comprehensive representation of the synthesized design in the CDFG format and accurately describes the hardware designs. Consequently, we use PIR to capture hardware design variations induced by pragmas. PIR CDFG is composed of nodes and edges, each labeled with specific types and attributes. A toylike code is illustrated in Fig. 1 (a) with no pragma applied, and its corresponding PIR generated through Vitis HLS [33] is illustrated in Fig. 1 (b). Node attributes include node types, bitwidths, and consts, while edge attributes include edge types and data flow directions. For example, the node add_ln7 of type 00, indexed at $18$, denotes an add operation with a bitwidth of $3$. The edge indexed at $40$ is of type $2$.
@@ V-C Comparison with the baselines
The two predictors in DID4444HLS share identical architecture, one for latency measured in clock cycles and the other for resource consumption measured in Sum of Normalized Resource Use (SNRU) [15, 42, 43]:
@@
$3$) SA-ML [16] is an annealing-based method. The predictive models are based on XGBoost [48]. We adopt Clang [^2] https://clang.llvm.org to compile the C++ code with pragmas into LLVM IRs. Then, the design features as input to the model are extracted from the IR CDFGs through Programl [20], including attributes such as the number of nodes of each type, the path length, and the pipelining weight. We set the initial temperature to $10$ with a temperature change rate of $0.1$ and $50$ runs at each temperature. Predictive models are trained with an initial set of $150$ designs randomly sampled and synthesized. Hyperparameter tuning is performed using Comet.ml with Bayesian optimization [^3] https://www.comet.com. We define a set of $30$ optimization weights evenly spaced between 00 and $1$ to approximate a Pareto front, and randomly sample $30$ designs as the starting point for each weight. Through simulated annealing, each design transitions to its neighbor by randomly selecting one of its pragma and changing to an adjacent value until the stopping condition, which is set to $500$ rejects or the temperature of $10^{-7}$. The candidates are synthesized after the final transitions.
@@ V-D Ablation study
Generative models are particularly relevant for learning conditional distributions, among which VAE, Generative Adversarial Network (GAN), and diffusion model have demonstrated state-of-the-art performance [37, 49, 50]. Therefore, we conducted an ablation study to compare their performance.
@@
Diffusion model stems from the idea that a continuous Gaussian diffusion process can be reversed with the same functional form as the forward process. The reverse process converges pure noise to samples in the target distribution. The diffusion process can be approximated with a discrete process given enough diffusion steps. The backbone of diffusion model is responsible for extracting meaningful features from the input data at various diffusion steps to iteratively remove noise. Fig. 9 shows the backbone structure we adapted from DiffWave [51], because the gated structure excels at filtering out critical information from the input data. Meanwhile, skip connections mitigate vanishing and exploding gradient problems and capture information at multiple levels of abstraction [51, 52, 53]. We set the number of diffusion steps as $50$ with a quadratic noise schedule from $\beta_{1}=10^{-4}$ to $\beta_{50}=0.5$ [52, 53].
@@ VI Discussion
GAN, while effective, is known to face challenges such as discriminator overfitting and training stability issues, especially in the cases of limited data availability [54, 55]. Due to the time-consuming nature of HLS, there is usually a scarcity of training data. The dataset is updated in each iteration and the model must be re-trained thereafter. Thus, the stability issues may accumulate throughout the optimization process and lead to failure in finding the directions along which the distributions converge toward the Pareto front. While diffusion model generates samples with high fidelity and diversity, the stochasticity introduced by multiple diffusion steps hinders the reparameterization trick during the gradient-based update of the distributions. LLVM IR combined with VAE performs far behind DID4444HLS, indicating the limitation of feature extraction from LLVM IR and the importance of using PIR to leverage post-HLS information.

### v3

@@ Abstract
High-level synthesis (HLS) has significantly advanced the automation of digital circuit design, yet the need for expertise and time in pragma tuning remains challenging. Existing solutions for the design space exploration (DSE) adopt either heuristic methods, lacking essential information for further optimization potential, or predictive models, missing sufficient generalization due to the time-consuming nature of HLS and the exponential growth of the design space. To address these challenges, we propose Deep Inverse Design for HLS (DID4444HLS), a novel approach that integrates graph neural networks and generative models. DID4HLS iteratively optimizes hardware designs aimed at compute-intensive algorithms by learning conditional distributions of design features from post-HLS data. Compared to four state-of-the-art DSE baselines, our method achieved an average improvement of $42.8\%$ on average distance to reference set (ADRS) compared to the best-performing baselines across six benchmarks, while demonstrating high robustness and efficiency. The code is available at https://github.com/PingChang818/DID4HLS.
Index Terms — deep learning, inverse design, GNN, HLS, ADRS
@@ I Introduction
Heuristic methods offer rapid implementations to yield improved designs [5]. However, their effectiveness is limited by the lack of a thorough understanding of the behaviors of HLS tools. We need to understand how operations are scheduled in order to estimate design performance, but scheduling information is only available after the HLS process [6]. Recently, graph neural networks (GNNs) have been employed for code analysis to predict synthesis results, demonstrating advantages over the preceding DSE methods [7, 8, 9, 10, 11]. Compilers transform algorithms in HLL into platform-independent intermediate representations (IRs) that encapsulate semantic information [12, 13, 14]. They are subsequently transformed into control and data flow graphs (CDFGs) for GNN training [15, 16]. However, three critical issues hinder the effectiveness of these methods:
$1$) The exponential increase in design space size and the time-consuming nature of HLS pose challenges in acquiring sufficient synthesis data for training high-generalization models [10].
@@
Gradient-based inverse design has been widely adopted to retrieve the nanophotonic device designs with desired quality of results (QoR) from large design spaces [18]. However, the inherent discrete nature of HLS design spaces keeps from applying the inverse design to DSE for HLS. Although several recent approaches approximated discrete spaces with continuous spaces [19], they are not feasible for the inverse design in HLS due to the fact that the pragma effect on the QoR is non-monotonic [8].
To address these challenges, we propose a novel method called Deep Inverse Design for HLS (DID4444HLS). DID4444HLS focuses on choosing the optimal pragma configurations of the compute-intensive algorithms to achieve designs that are close to Pareto points. We construct the design space with the pragmas of pipelining, loop unrolling, array partitioning, and function inlining. The uniqueness of DID4444HLS lies in its strategy to work on a distribution that samples pragma configurations, with constraints that prevent redundant design samples according to the strategies of HLS tools, such as interactions between pipelining and loop unrolling. Initially, DID4444HLS uses a random pragma distribution to explore a wide range of design possibilities. It then iteratively synthesizes batches of designs generated from the updated pragma distributions, resulting in the IRs that are generally represented in CDFG and crucial in deriving their final hardware designs [20]. In this work, we call these IRs as post-HLS IRs (PIRs). These PIRs, along with data on their objectives, are collected into a dataset which is used to train the objective predictors based on Graph Attention Network v2222 (GATv2222) [21]. Furthermore, DID4444HLS adopts Variational Autoencoder (VAE) to learn the distribution of design features that determine design performance. They are extracted from the PIRs by the GATv2222 models. This training is conditioned on the pragma distributions. Through this process, DID4444HLS effectively converges the pragma distribution to approximate Pareto front [3].
@@ II-A Post-HLS Intermediate Representation
The initial HLS phase parses algorithms in HLL into virtual instruction sets called intermediate representations. Subsequent phases entail specifying the order and timing of these instructions under predefined constraints to generate PIRs [20]. PIR is a comprehensive representation of the synthesized design in the CDFG format and accurately describes the hardware designs. Consequently, we use PIR to capture hardware design variations induced by pragmas. PIR CDFG is composed of nodes and edges, each labeled with specific types and attributes. A toylike code is illustrated in Fig. 1 (a) with no pragma applied, and its corresponding PIR generated through Vitis HLS [24] is illustrated in Fig. 1 (b). Node attributes include node types, bitwidths, and consts, while edge attributes include edge types and data flow directions. For example, the node add_ln7 of type 00, indexed at $18$, denotes an add operation with a bitwidth of $3$. The edge indexed at $40$ is of type $2$.
@@ V-C Comparison with the baselines
The two predictors in DID4444HLS share identical architecture, one for latency measured in clock cycles and the other for resource consumption measured in Sum of Normalized Resource Use (SNRU) [10]:
@@
$3$) SA-ML [11] is an annealing-based method. The predictive models are based on XGBoost. We adopt Clang [^2] https://clang.llvm.org to compile the C++ code with pragmas into LLVM IRs. Then, the design features as input to the model are extracted from the IR CDFGs through Programl [15], including attributes such as the number of nodes of each type, the path length, and the pipelining weight. We set the initial temperature to $10$ with a temperature change rate of $0.1$ and $50$ runs at each temperature. Predictive models are trained with an initial set of $150$ designs randomly sampled and synthesized. Hyperparameter tuning is performed using Comet.ml with Bayesian optimization [^3] https://www.comet.com. We define a set of $30$ optimization weights evenly spaced between 00 and $1$ to approximate a Pareto front, and randomly sample $30$ designs as the starting point for each weight. Through simulated annealing, each design transitions to its neighbor by randomly selecting one of its pragma and changing to an adjacent value until the stopping condition, which is set to $500$ rejects or the temperature of $10^{-7}$. The candidates are synthesized after the final transitions.
@@ V-D Ablation study
Generative models are particularly relevant for learning conditional distributions, among which VAE, Generative Adversarial Network (GAN), and diffusion model have demonstrated state-of-the-art performance [28]. Therefore, we conducted an ablation study to compare their performance.
@@
Diffusion model stems from the idea that a continuous Gaussian diffusion process can be reversed with the same functional form as the forward process. The reverse process converges pure noise to samples in the target distribution. The diffusion process can be approximated with a discrete process given enough diffusion steps. The backbone of diffusion model is responsible for extracting meaningful features from the input data at various diffusion steps to iteratively remove noise. Fig. 9 shows the backbone structure we adapted from DiffWave [37], because the gated structure excels at filtering out critical information from the input data. Meanwhile, skip connections mitigate vanishing and exploding gradient problems and capture information at multiple levels of abstraction [37, 38]. We set the number of diffusion steps as $50$ with a quadratic noise schedule from $\beta_{1}=10^{-4}$ to $\beta_{50}=0.5$ [38].
@@ VI Discussion
GAN, while effective, is known to face challenges such as discriminator overfitting and training stability issues, especially in the cases of limited data availability [39]. Due to the time-consuming nature of HLS, there is usually a scarcity of training data. The dataset is updated in each iteration and the model must be re-trained thereafter. Thus, the stability issues may accumulate throughout the optimization process and lead to failure in finding the directions along which the distributions converge toward the Pareto front. While diffusion model generates samples with high fidelity and diversity, the stochasticity introduced by multiple diffusion steps hinders the reparameterization trick during the gradient-based update of the distributions. LLVM IR combined with VAE performs far behind DID4444HLS, indicating the limitation of feature extraction from LLVM IR and the importance of using PIR to leverage post-HLS information.